# Teaching Small Language Models to Learn Logic through Meta-Learning

### v2

@@ 1 Introduction
With the advent of increasingly capable large language models (LLMs), logical reasoning has become a central domain for evaluating and comparing these systems (Huang and Chang, 2023; Mondorf and Plank, 2024; Liu et al., 2025). However, language models’ ability to perform rigorous logical reasoning remains an open question. Even with extensive scientific, mathematical, and programming data, LLMs still fail to generalize beyond seen inference patterns (Clark et al., 2021; Saparov et al., 2023; Mirzadeh et al., 2025; Gulati et al., 2024; Huang et al., 2025; Shi et al., 2023; Yoran et al., 2024). These challenges relate to broader debates surrounding generalization versus memorization in LLMs (Balloccu et al., 2024; Singh et al., 2024).
@@
Few-shot meta-learning approaches (Irie and Lake, 2024) have emerged as promising methods for inducing OOD generalization and rapid domain adaptation in LLMs. Specifically, this class of methods has proven effective in few-shot task generalization (Min et al., 2022; Chen et al., 2022), systematic generalization (Lake and Baroni, 2023), and mitigating catastrophic forgetting (Irie et al., 2025). In this work, we propose for the first time to apply few-shot meta-learning (ML) to enhance the ability of LLMs to learn logic.
As illustrated in Figure 1, we evaluate the effectiveness of this approach using a logical reasoning task grounded in syllogistic logic (Smiley, 1973; Vargas Guzmán et al., 2024) and frame it as a premise selection task. This is a minimal but still interesting and non-trivial fragment of logic consisting of a few clear logical rules. As such it allows us to directly address our research question: Does ML enhance LLM’s ability to learn abstract logical patterns underlying reasoning and generalization? Each problem presents a knowledge base of atomic logical statements. Models are tasked with identifying the minimal subset of premises that logically entail a given test hypothesis. This premise selection task captures a core aspect of deductive reasoning: determining which known facts are necessary and sufficient to justify a conclusion. We apply ML to small LMs from the Qwen-2.5 family (Qwen Team, 2025), ranging from 1.5B to 7B parameters. Specifically, we assess the generalization capabilities induced by ML, such as systematically performing inferences over unseen sets of premises, as well as over more complex (longer) or simpler (shorter) sets of premises than those encountered during training. [^1] Our code and data are available at: https://github.com/leobertolazzi/meta-learning-logic.git.
Our main contributions are as follows: We introduce a new synthetic dataset based on syllogistic logic to directly study logical reasoning generalization in LLMs. We show that ML enables small LMs to better generalize in OOD syllogistic reasoning problems with particularly strong performance in smaller models and low-data regimes. We demonstrate that small LMs fine-tuned with ML can outperform state-of-the-art LLMs such as GPT-4o and o3-mini, on our syllogistic premise selection task.
@@ 4.1 Models and Evaluation
Fine-tuning
Qwen-2.5 1.5B ML 93.11 ± 0.61 94.28 ± 0.61 91.76 ± 0.27
@@ 6 Error Analysis
Beyond simply measuring the accuracy of ML and the baseline, we additionally focus on two main types of errors models make when evaluated on unseen lengths. First, among all errors, we consider the proportion of non-minimal valid set of premises (NVM). This means that the correct minimal set was generated by the model, but together with unnecessary premises; for this case, we also measure how many unnecessary premises, on average, the models generate. Alternatively, models may fail to provide the complete $A$-chain within the correct minimal set of premises, meaning that at least one necessary $A$ premise is missing (MAP); here, we also track the average number of missing necessary $A$-formulas in erroneous answers. NVM and MAP are mutually exclusive. Furthermore, we consider an additional type of error that can occur simultaneously with either NVM or MAP: models may hallucinate premises —referred to as hallucinated premises (HP) —and output a formula that is not contained in the $\mathcal{KB}$.
@@ 7.1 LLMs’ Logical Capabilities
LLMs often struggle with OOD generalization (Clark et al., 2021; Saparov et al., 2023; Vargas Guzmán et al., 2024), multi-step inference (Creswell et al., 2023), and consistency across formal reasoning patterns (Parmar et al., 2024; Hong et al., 2024). Neuro-symbolic methods address these gaps by integrating logic modules or symbolic solvers, improving both performance and interpretability (Pan et al., 2023; Olausson et al., 2023; Kambhampati et al., 2024). In a different direction, LRMs have shown strong gains in reasoning and planning tasks (Xu et al., 2025). Our proposed meta-learning approach offers a complementary alternative by enabling LLMs to systematically learn logical relations.
@@ 7.2 Meta-learning
Meta-learning enables models to rapidly adapt to new tasks by leveraging prior experiences across tasks (Thrun and Pratt, 1998; Hospedales et al., 2022). Foundational approaches include memory-augmented neural networks (Santoro et al., 2016), prototypical networks (Snell et al., 2017), and model-agnostic meta-learning (MAML) (Finn et al., 2017). In the context of LLMs, meta-learning has been explored through techniques such as meta-in-context learning (Coda-Forno et al., 2023), in-context tuning (Chen et al., 2022), and MetaICL (Min et al., 2022), which either train for or exploit the in-context learning abilities of models to adapt to new tasks using few-shot examples. Our proposed method draws inspiration from the MSL framework (Irie and Lake, 2024), which we adapt and extend to solve the logical premise selection task.
@@ D.3 Generalization with Limited Data
D.4 Providing Study Examples to the Baseline at Inference Time
@@
D.5 Additional Error Analysis
@@
Core
Qwen-2.5 1.5B $\mathrm{Baseline}_{-S}$ 85.56 ± 1.24 91.42 ± 0.82 80.56 ± 1.78

### v3

@@ 1 Introduction
Historically, a central debate in AI has focused on whether neural networks, which are often characterized as associative statistical systems, can capture abstract rules and learn to apply them systematically (Fodor and Pylyshyn, 1988; Smolensky, 1990; Lake and Baroni, 2018; Hupkes et al., 2020; Kim and Linzen, 2020; McCurdy et al., 2024).
In recent years, this debate has resurfaced when investigating Large Language Models’ (LLMs) reasoning capabilities (Huang and Chang, 2023; Mondorf and Plank, 2024; Liu et al., 2025). Their ability to perform rigorous logical reasoning is often contested, raising the question of whether they achieve true generalization or merely rely on memorized patterns from their training data (Balloccu et al., 2024; Singh et al., 2024). Even with extensive scientific, mathematical, and programming data, LLMs still fail to generalize beyond seen inference patterns (Clark et al., 2021; Saparov et al., 2023; Mirzadeh et al., 2025; Gulati et al., 2024; Huang et al., 2025; Shi et al., 2023; Yoran et al., 2024).
@@
In this context, few-shot meta-learning (ML) approaches (Irie and Lake, 2024) have emerged as promising methods for inducing out-of-distribution (OOD) generalization and rapid domain adaptation in LLMs. Specifically, this class of methods has proven effective in few-shot task generalization (Min et al., 2022; Chen et al., 2022), mitigating catastrophic forgetting (Irie et al., 2025), and inducing systematic generalization in rule induction tasks (Lake and Baroni, 2023).
However, it remains unclear whether these findings would translate to the domain of deductive reasoning and logic, where models are tasked to learn and apply more complex rules than what has been studied in previous research. Can few-shot meta-learning (ML) enable LLMs to learn the abstract logical patterns underlying reasoning and generalization within a formal system? To answer this foundational question, we propose, for the first time, to apply ML to enhance the ability of LLMs to learn logic within a rigorously defined and controlled fragment of natural language (Pratt-Hartmann, 2004). As illustrated in Figure 1, we evaluate the effectiveness of ML using a logical reasoning task grounded in syllogistic logic (Smiley, 1973; Vargas Guzmán et al., 2024) and frame it as a premise selection task. This is a minimal but still interesting and non-trivial fragment of logic consisting of a few clear logical rules. Each problem presents a knowledge base of atomic logical statements. Models are tasked with identifying the minimal subset of premises that logically entail a given test hypothesis. This premise selection task captures a core aspect of deductive reasoning: determining which known facts are necessary and sufficient to justify a conclusion. We apply ML to small LMs from the Qwen-2.5 family (Qwen Team, 2025), ranging from 1.5B to 7B parameters. Specifically, we assess the generalization capabilities induced by ML, such as systematically performing inferences over unseen sets of premises, as well as over more complex (longer) or simpler (shorter) sets of premises than those encountered during training. [^1] Our code and data are available at: https://github.com/leobertolazzi/meta-learning-logic.git.
Our main contributions are as follows:
  • We introduce a new synthetic dataset based on syllogistic logic to directly study logical reasoning generalization in LLMs within a controlled setting.
  • We show that ML enables small LMs to better generalize in OOD syllogistic reasoning problems with particularly strong performance in smaller models and low-data regimes.
  • We demonstrate that small LMs fine-tuned with ML can outperform state-of-the-art LLMs such as GPT-4o and o3-mini, on our syllogistic premise selection task.
@@ 4.1 Models and Evaluation
Fine-tuning Qwen-2.5 1.5B ML 93.11 ± 0.61 94.28 ± 0.61 91.76 ± 0.27
@@ 5.1 Core Generalization
For completeness, we also include the results of Qwen-2.5 models in a few-shot prompting setting in Appendix D.4. These results demonstrate that the base model, prior to fine-tuning, is indeed largely incapable of performing the task.
@@ 6 Error Analysis
Beyond simply measuring the accuracy of ML and the baseline, we additionally focus on two main types of errors models make when evaluated on unseen lengths. First, among all errors, we consider the proportion of non-minimal valid set of premises (NVM). This means that the correct minimal set was generated by the model, but together with unnecessary premises; for this case, we also measure how many unnecessary premises, on average, the models generate. Note that a very low number of unnecessary premises indicates that a model is indeed close to the correct solution, whereas including many unnecessary premises, with the extreme case of outputting the full knowledge base, would indicate that models learned to exploit a trivial solution instead of internalizing the abstract inference pattern beyond the length seen during training. Alternatively, models may fail to provide the complete $A$-chain within the correct minimal set of premises, meaning that at least one necessary $A$ premise is missing (MAP); here, we also track the average number of missing necessary $A$-formulas in erroneous answers. NVM and MAP are mutually exclusive. Furthermore, we consider an additional type of error that can occur simultaneously with either NVM or MAP: models may hallucinate premises —referred to as hallucinated premises (HP) —and output a formula that is not contained in the $\mathcal{KB}$.
@@
Overall, the fact that both ML models and the baseline have a higher percentage of NVM errors when generalizing to shorter inferences and, conversely, a higher percentage of MAP errors when generalizing to longer inferences shows that these models still exhibit some bias towards the lengths seen during training, reflected in the correlation between training length and error types at test time.
@@ 7.1 LLMs’ Logical Capabilities
LLMs often struggle with OOD logical generalization (Clark et al., 2021; Saparov et al., 2023; Vargas Guzmán et al., 2024), multi-step inference (Creswell et al., 2023), and consistency across formal reasoning patterns (Parmar et al., 2024; Hong et al., 2024). Neuro-symbolic methods address these gaps by integrating logic modules or symbolic solvers, improving both performance and interpretability (Pan et al., 2023; Olausson et al., 2023; Kambhampati et al., 2024). In a different direction, LRMs have shown strong gains in reasoning and planning tasks (Xu et al., 2025). Our proposed meta-learning approach offers a complementary alternative by enabling LLMs to systematically learn logical relations without relying on symbolic modules.
Many benchmarks, such as LogiQA (Liu et al., 2020), ProofWriter (Tafjord et al., 2021), FOLIO (Han et al., 2024), and ProntoQA (Saparov et al., 2023), have been proposed to measure logical reasoning in LLMs. However, we relied on syllogistic logic since existing datasets either do not control for specific types of inferences or only include a limited set of very basic inference rules, e.g., modus ponens. Moreover, most of these datasets do not allow us to vary the complexity (length) of inferences or the set of premises forming knowledge bases from which inferences can be drawn.
@@ 7.2 Meta-learning
Meta-learning enables models to rapidly adapt to new tasks by leveraging prior experiences across tasks (Thrun and Pratt, 1998; Hospedales et al., 2022). Foundational approaches include memory-augmented neural networks (Santoro et al., 2016), prototypical networks (Snell et al., 2017), and model-agnostic meta-learning (MAML) (Finn et al., 2017). In the context of LLMs, meta-learning has been explored through techniques such as meta-in-context learning (Coda-Forno et al., 2023), in-context tuning (Chen et al., 2022), and MetaICL (Min et al., 2022), which either train for or exploit the in-context learning abilities of models to adapt to new tasks using few-shot examples. Our proposed method draws inspiration from the MSL framework (Irie and Lake, 2024), which we adapt and extend to solve the logical premise selection task. Within the meta-learning literature, Lake and Baroni (2023) is most closely related to our work. While they explore how meta-learning induces systematic generalization by composing simple primitive functions with few arguments, we instead focus on fixed, more complex abstract rules rooted in syllogistic logic, featuring a varying number of premises from which an inference can be drawn.
@@ 9 Limitations
10 Acknowledgement
This work was partially funded by the National Science Centre, Poland under the OPUS call [grant 2020/37/B/HS1/04220]. We gratefully acknowledge Polish high-performance computing infrastructure PLGrid (HPC Center: ACK Cyfronet AGH) for providing computer facilities and support within computational grant no. PLG/2025/018321.
@@ D.3 Generalization with Limited Data
D.4 Comparison with Prompted Base Model
To better contextualize the base Qwen-2.5 models’ capabilities prior to any fine-tuning, Table 9 presents the performance of Qwen-2.5 models using few-shot prompting with greedy decoding compared with the performance of the ML and baseline models previously shown in the main paper. The few-shot examples are the same as the one provided to the ML and $\mathrm{Baseline}_{+S}$ models at test time. This comparison allows us to verify that the base models cannot perform premise selection effectively through in-context learning alone, thereby justifying the need for studying how different fine-tuning approaches can affect how the model learns logical inferences.
The results show dramatically poor performance across all model sizes: even the largest 7B model achieves only 11.14% accuracy overall, while the smallest 1.5B model reaches merely 1.09%. Although performance improves modestly with model scale, all results remain far below both the fine-tuned baseline and meta-learning approaches, which exceed 85% accuracy even at the smallest size. Notably, even in this low-performance regime, the pattern observed in fine-tuned models persists: longer inferences remain more difficult than shorter ones across all model sizes (e.g., for the 7B model, accuracy drops from 13.89% on short inferences to 10.49% on long ones), further supporting our hypothesis that selecting all premises from a larger minimal set is more challenging.
Model Method All Short Long
Qwen-2.5 1.5B ML 93.11 ± 0.61 94.28 ± 0.61 91.76 ± 0.27
$\mathrm{Baseline}_{-S}$ 85.56 ± 1.24 91.42 ± 0.82 80.56 ± 1.78
Few-shot Prompting 1.09 ± 0.00 2.00 ± 0.00 0.51 ± 0.00
Qwen-2.5 3B ML 96.16 ± 0.44 96.24 ± 0.56 95.55 ± 0.43
$\mathrm{Baseline}_{-S}$ 93.03 ± 1.15 95.34 ± 1.18 90.92 ± 1.27
Few-shot Prompting 4.98 ± 0.00 8.20 ± 0.00 3.11 ± 0.00
Qwen-2.5 7B ML 98.13 ± 0.98 98.26 ± 0.82 97.69 ± 1.40
$\mathrm{Baseline}_{-S}$ 95.76 ± 1.10 97.27 ± 1.22 94.13 ± 0.90
Few-shot Prompting 11.14 ± 0.00 13.89 ± 0.00 10.49 ± 0.00
Table 9: Comparison with prompted Qwen-2.5 models on core generalization. Accuracy (mean ± std) on test inferences across all type‐length combinations (All), plus breakdown into the five shortest (Short) and longest (Long) inferences for each of the seven types of inference. The table includes the results obtained by a few-shot prompted Qwen-2.5 models using greedy decoding.
D.5 Providing Study Examples to the Baseline at Inference Time
@@
D.6 Additional Error Analysis
@@
Core Qwen-2.5 1.5B $\mathrm{Baseline}_{-S}$ 85.56 ± 1.24 91.42 ± 0.82 80.56 ± 1.78